OpenStudy (anonymous):

How wld I solve this or it doesnt matter? ||u||^2

OpenStudy (anonymous):

like shld i first find the magnitude of one and then square it?

OpenStudy (anonymous):

sorry its worded funny but i cant think of any other way

OpenStudy (turingtest):

\[u=(a,b,c)\]\[||u||=\sqrt{a^2+b^2+c^2}\]so\[||u||^2=a^2+b^2+c^2\]so it's actually a little less work than finding the magnitude usually

OpenStudy (anonymous):

kk i see thanks